How do i slove d= 8.3 m to get an area of a circle
Answer by rothauserc(4718)

I assume d = 8.3 is the diameter of the circle
therefore the radius of the circle is 8.3/2 = 4.15 m
the area of the circle equals pi * radius^2
Area = 22/7 * 4.15^2 = 54.10 square m